OpenStudy (anonymous):

how to find the limit as n->infinity of: (1/3)*(e^-nx)*n^2

OpenStudy (freckles):

are you allowed to use l'hospital?

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}\frac{e^{-nx} n^2}{3}=\lim_{n \rightarrow \infty}\frac{n^2}{3 e^{nx}}\]

OpenStudy (freckles):

since you have inf/inf apply l'hospital

OpenStudy (anonymous):

ah of course! Thank you.

OpenStudy (freckles):

np